OpenStudy (bry856):

If (2 − 3i) + (x + yi) = 6, what is x + yi?

OpenStudy (bry856):

@Preetha

OpenStudy (johnweldon1993):

Are there options on this one?

OpenStudy (bry856):

yeah hold on ill screenshot it

OpenStudy (johnweldon1993):

Okay, just wanted to make sure it wasnt more complicated than it is \[\large (2 - 3i) + (x + yi) = 6\] All we need to do is rearrange this to solve for (x + yi) \[\large x + yi = 6 - (2 - 3i)\] \[\large x + yi = 4 - 3i\]

OpenStudy (bry856):

thank you so much!!

OpenStudy (johnweldon1993):

Of course :)